# Data-based control of Logical Networks

### v1

@@ A Boolean control network (BCN) is a logical system described by the following equations
and the associated digraph $\mathcal{D}(L_{tot})\triangleq(\mathcal{V},\mathcal{E}_{tot})$. The state ${\bf x}_{f}=\delta_{N}^{i}$ is reachable from ${\bf x}_{0}=\delta_{N}^{j}$ if and only if there exists $\tau\in\mathbb{Z}_{+}$ such that $[L_{tot}^{\tau}]_{ij}=1$, or equivalently, there is a path of length $\tau$ from $j$ to $i$ in $\mathcal{D}(L_{tot})$. Consequently, a BCN is reachable if and only if the matrix $L_{tot}$ is irreducible, or equivalently $\mathcal{D}(L_{tot})$ is strongly connected. In the following, we will denote by $\mathcal{R}({\bf x}_{0})$ the set of states reachable from ${\bf x}_{0}$.
@@
Note that ${\bf x}_{e}=\delta^{j}_{N}$ is an equilibrium point of a BN or a BCN if and only if the node $j$ has a self-loop in ${\mathcal{D}}(L)$ or ${\mathcal{D}}(L_{tot})$, respectively.
@@
Cheng & Liu (2009); Cheng et al. (2011); Fornasini & Valcher (2013b) A BCN (4a) is stabilizable to the state ${\bf x}_{e}\in\mathcal{L}_{N}$ if for every ${\bf x}(0)\in\mathcal{L}_{N}$ there exist $\tau\in\mathbb{Z}_{+}$ and ${\bf u}(t),t\in\mathbb{Z}_{+}$, that leads to ${\bf x}(t)={\bf x}_{e},\forall t\geq\tau$.
Stabilizability can be characterized as follows.
Proposition 4.
Cheng et al. (2011); Fornasini & Valcher (2013b); Li et al. (2013) A BCN (4a) is stabilizable to ${\bf x}_{e}\in\mathcal{L}_{N}$ if and only if the following conditions hold:
  • (i)
    ${\bf x}_{e}$ is an equilibrium point of the BCN corresponding to some input ${\bf u}=\delta^{i}_{M}$, $i\in[1,M]$;
  • (ii)
    ${\bf x}_{e}$ is reachable from every state ${\bf x}(0)$, i.e., ${\bf x}_{e}\in\bigcap_{{\bf x}(0)\in\mathcal{L}_{N}}\mathcal{R}({\bf x}(0))$ (i.e., ${\bf x}_{e}$ is globally reachable).
@@
If we define the basin of attraction of a state ${\bf x}_{e}\in{\mathcal{L}}_{N}$ as the set of states ${\bf x}\in{\mathcal{L}}_{N}$ from which it is possible to reach ${\bf x}_{e}$ and denote it by ${\mathcal{S}}({\bf x}_{e})$, then condition (ii) in Proposition 4 is equivalent to ${\mathcal{S}}({\bf x}_{e})={\mathcal{L}}_{N}$.
By Proposition 4, in order to identify the states to which system (4a) can be stabilized, we first need to determine the set $\mathcal{X}_{e}$ of the equilibrium points of the BCN (corresponding to the various inputs), namely the nodes of $\mathcal{D}(L_{tot})$ with a self-loop. Then, we need to identify the states in $\mathcal{X}_{e}$ that are reachable from every other state, namely the states in $\mathcal{X}_{e}$ whose associated node in $\mathcal{D}(L_{tot})$ is globally reachable. We will henceforth interchangeably write that a node (in the digraph) or a state (of the BCN) are globally reachable.
It has been shown in Cheng et al. (2011); Fornasini & Valcher (2013b); Li et al. (2013) that if a BCN is stabilizable to some ${\bf x}_{e}$, then it is stabilizable by means of a state feedback law: ${\bf u}(t)=K{\bf x}(t),\exists K\in{\mathcal{L}}_{M\times N}.$ When so the BCN becomes a BN:
${\bf x}(t+1)=L\ltimes K\ltimes{\bf x}(t)\ltimes{\bf x}(t)=\tilde{L}(K){\bf x}(t),$ (8)
for some $\tilde{L}(K)\in{\mathcal{L}}_{N\times N}.$
3 Data-driven stabilization
In this section we restrict our attention to the state equation (4a) of the BCN. We assume to have performed an offline experiment during which we have collected some state/input data from the BCN (4a) on a finite time interval $[0,T],T\in\mathbb{Z}_{+}$111 Instead of a single experiment to collect data, we can perform several experiments, say $r$, on some finite time intervals $[0,T_{i}],i\in[1,r]$. It suffices to arrange data consistently with the way it is illustrated in the case of a single experiment., with $T$ sufficiently large (in a sense that will be clarified in the sequel), and we define the vector sequences ${\bf x}_{d}\triangleq\{{\bf x}_{d}(t)\}_{t=0}^{T}$, and ${\bf u}_{d}\triangleq\{{\bf u}_{d}(t)\}_{t=0}^{T-1}$. We rearrange the collected data into the following matrices:
$X_{p}$$\triangleq$ $\begin{bmatrix}{\bf x}_{d}(0)&{\bf x}_{d}(1)&\dots&{\bf x}_{d}(T-1)\end{bmatrix}\in\mathcal{L}_{N\times T},$ (9)
$X_{f}$$\triangleq$ $\begin{bmatrix}{\bf x}_{d}(1)&{\bf x}_{d}(2)&\dots&{\bf x}_{d}(T)\end{bmatrix}\in\mathcal{L}_{N\times T},$ (10)
$U_{p}$$\triangleq$ $\begin{bmatrix}{\bf u}_{d}(0)&{\bf u}_{d}(1)&\dots&{\bf u}_{d}(T-1)\end{bmatrix}\in\mathcal{L}_{M\times T},$ (11)
where the subscripts $p$ and $f$ stand for past and future, respectively.
As a BCN (4a) bijectively corresponds to its describing matrix, we define the set of Boolean control networks (4a) compatible with the data $({\bf x}_{d}$, ${\bf u}_{d})$ as
${\mathcal{B}}_{d}$$\triangleq$ $\{\tilde{L}\in\mathcal{L}_{N\times NM}:$
${\bf x}_{d}(t+1)=\tilde{L}\ltimes{\bf u}_{d}(t)\ltimes{\bf x}_{d}(t),\forall t\in[0,T-1]\}.$
Obviously, the matrix $L$ of the BCN that has generated the data belongs to ${\mathcal{B}}_{d}$. However, ${\mathcal{B}}_{d}$ may also contain other matrices if the data do not allow to uniquely determine $L$. To this end, we introduce the notion of informativity for identifiability.
Definition 5.
We say that the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for identifiability of the BCN (4a) if ${\mathcal{B}}_{d}\equiv\{L\}$.
We can provide the following characterization of informativity for identifiability.
Lemma 6.
The data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for identifiability of the BCN (4a) if and only if the logical matrix
$\left[\begin{array}[]{c|c|c}{\rm col}_{1}(U_{p})\ltimes{\rm col}_{1}(X_{p})&\dots&{\rm col}_{T}(U_{p})\ltimes{\rm col}_{T}(X_{p})\end{array}\right]$ (13)
is of full row rank.
Proof.
A BCN (4a) is uniquely identified from the data matrices $X_{p}$, $X_{f}$ and $U_{p}$ if and only if these latter allow to deduce each column of the matrix $L$ in (4a), or, equivalently, are representative of every possible transition that can occur in the network (from all possible state vectors and corresponding to all possible inputs). By taking into account that the $k$-th column of the matrix $X_{f}$ represents the successor state of the state given by the $k$-th column of $X_{p}$, when the applied input is the one corresponding to the $k$-th column of $U_{p}$, it follows that we know from data every possible transition of the BCN if and only if for every $(i,j)\in[1,M]\times[1,N]$ there exists $k\in[1,T]$ such that
$\begin{bmatrix}X_{p}\cr U_{p}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{M}^{i}\end{bmatrix},$ (14)
which, by resorting to the semi-tensor product, is equivalent to ${\rm col}_{k}(U_{p})\ltimes{\rm col}_{k}(X_{p})=\delta_{M}^{i}\ltimes\delta_{N}^{j}.$ But, saying that for every $(i,j)\in[1,M]\times[1,N]$ condition (14) holds for some $k\in[1,T]$ amounts to saying that the matrix (13) contains all vectors of $\mathcal{L}_{NM}$, i.e., is of full row rank. ∎
Remark 7.
If (14) holds, then $X_{f}\delta_{T}^{k}={\rm col}_{j}(L_{i}).$
To make the following analysis meaningful, from now on we assume that the collected data are not informative for identifiability of the considered BCN, namely ${\mathcal{B}}_{d}\supsetneq\{L\}$, and we investigate under what conditions it is still possible to use them to solve some control problems.
We now formalize the data-driven state feedback stabilization problem.
Problem 1.
Given the data $({\bf x}_{d}$, ${\bf u}_{d})$, determine (if it exists) a state feedback matrix $K\in\mathcal{L}_{M\times N}$ such that ${\bf u}(t)=K{\bf x}(t)$ stabilizes all the BCNs compatible with the data to some state ${\bf x}_{e}\in\mathcal{L}_{N}$.
When Problem 1 is solvable, we will say that the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for stabilization to ${\bf x}_{e}\in\mathcal{L}_{N}$ by state feedback.
By Proposition 4, it is possible to stabilize a BCN to some state ${\bf x}_{e}\in{\mathcal{L}}_{N}$ only if such a state is an equilibrium point of the BCN. Therefore Problem 1 is solvable only if ${\bf x}_{e}$ is an equilibrium point for all possible BCNs compatible with the data, namely all the BCNs in ${\mathcal{B}}_{d}$. By Definition 2, a state ${\bf x}_{e}$ is an equilibrium point of a BCN (4a) if there exists $i\in[1,M]$ such that for ${\bf u}=\delta_{M}^{i}$ the successor of ${\bf x}_{e}$ is ${\bf x}_{e}$ itself. Therefore, in order to identify the equilibrium points compatible with the collected data, we can stack the matrices $X_{p}$ and $X_{f}$ one above the other, and check in which columns the logical vector in the upper matrix is the same as the logical vector in the lower matrix. Such a logical vector is an equilibrium point, since the column of $X_{f}$ represents the successor of the corresponding column in $X_{p}$ (for some input). More formally, we can apply the following algorithm.
Algorithm 1: Equilibrium points of the BCNs in ${\mathcal{B}}_{d}$
Input: The data matrices $X_{p}$ and $X_{f}$.
Output: The set $\mathcal{X}_{e}^{d}$ of all equilibrium points ${\bf x}_{e}$ compatible with the data.
Initialization:
Set $\bar{X}=\mathbbold{0}_{N\times T}$.
Procedure:
for $(j,k)\in[1,N]\times[1,T]$, do
  • $[\bar{X}]_{jk}=[X_{p}]_{j,k}[X_{f}]_{j,k}$
end for
Set:
  • $\{(j_{1},k_{1}),(j_{2},k_{2}),\dots,(j_{n_{e}},k_{n_{e}})\}$ = nonzero$(\bar{X})$
  • $\mathcal{X}_{e}^{d}=\{\delta_{N}^{j_{1}},\delta_{N}^{j_{2}},\dots,\delta_{N}^{j_{n_{e}}}\}$.
Conclusion: Return $\mathcal{X}_{e}^{d}$. ∎
Note that Algorithm 1 first calculates the Hadamard product $\bar{X}=X_{p}\odot X_{f}$ of the logical matrices $X_{p}$ and $X_{f}$, and then identifies the $n_{e}\geq 0$ positions of the nonzero entries of $\bar{X}$: the first (i.e., row) indices of these positions are the indices of the logical vectors that represent equilibrium points of all the BCNs in ${\mathcal{B}}_{d}$.
Once we have determined the equilibria in $\mathcal{X}_{e}^{d}$, we can identify the input(s) to which they correspond. Indeed, if ${\bf x}_{e}=\delta_{N}^{j}\in\mathcal{X}_{e}^{d}$, it means that $\exists k\in[1,T]$ such that
$\bar{X}\delta_{T}^{k}=\delta_{N}^{j}\quad\Leftrightarrow\quad\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{j}\end{bmatrix}.$
Therefore, ${\bf u}_{j}\triangleq U_{p}\delta_{T}^{k}=\delta_{M}^{i},\exists i\in[1,M]$, is the/a corresponding input. Once we know $\mathcal{X}_{e}^{d}$, we can verify if the state ${\bf x}_{e}$ to which we want to stabilize the BCNs in ${\mathcal{B}}_{d}$ belongs to $\mathcal{X}_{e}^{d}$. If not, the problem is not solvable, while if ${\bf x}_{e}\in\mathcal{X}_{e}^{d}$ we need to verify (using only data) that condition (ii) of Proposition 4 holds, namely that ${\bf x}_{e}$ is globally reachable, in all the networks in ${\mathcal{B}}_{d}$. To this end, we introduce the notion of informativity for reachability of a generic $\bar{\bf x}\in{\mathcal{L}}_{N}$.
Definition 8.
Given $\bar{\bf x}\in\mathcal{L}_{N}$, we say that the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for reachability of $\bar{\bf x}$ if $\bar{\bf x}$ is globally reachable in all the BCNs in ${\mathcal{B}}_{d}$.
To test informativity for reachability, we propose the following algorithm.
Algorithm 2: Reachability of $\bar{\bf x}\in\mathcal{L}_{N}$
Input: - The data matrices $X_{p},X_{f},$ and $U_{p}$;
           - $\bar{\bf x}=\delta_{N}^{j}\in\mathcal{L}_{N}$.
Output: - Is $\bar{\bf x}$ globally reachable for every BCN in ${\mathcal{B}}_{d}$?
               Yes/No.
             - The basin of attraction $\mathcal{S}(\bar{\bf x})$ of $\bar{\bf x}$.
Initialization:
Set $d=0$, ${\mathcal{S}}_{0}=\{\delta_{N}^{j}\}$, $({\bf u}_{1}^{\bar{\bf x}},{\bf u}_{2}^{\bar{\bf x}},\dots,{\bf u}_{N}^{\bar{\bf x}})=(\mathbbold{0},\mathbbold{0},\dots,\mathbbold{0})$.
Iterative procedure:
if ${\mathcal{S}}_{d}\neq\emptyset$, then
  • $d\leftarrow d+1$
  • ${\mathcal{S}}_{d}=\emptyset$
  • for $k\in[1,T]$, do
            if
    $\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{r}\cr\delta_{N}^{q}\end{bmatrix},\qquad\exists\ \delta^{q}_{N}\in{\mathcal{S}}_{d-1}$
    and $\delta^{r}_{N}\not\in\cup_{i=1}^{d}{\mathcal{S}}_{i}$, then
                   ${\mathcal{S}}_{d}\leftarrow{\mathcal{S}}_{d}\cup\{\delta^{r}_{N}\}$ and ${\bf u}_{r}^{\bar{\bf x}}\triangleq U_{p}\delta_{T}^{k}$
            end if
    end for
  • Go back to the Iterative procedure.
else set $\mathcal{S}(\bar{\bf x})=\bigcup_{i=0}^{d-1}\mathcal{S}_{i}$ and go to the Conclusion
end if
Conclusion: if $\mathcal{S}(\bar{\bf x})=\mathcal{L}_{N}$, then return (‘Yes’, $\mathcal{S}(\bar{\bf x})$)
                 else return (‘No’, $\mathcal{S}(\bar{\bf x})$)
                 end if
Note that the Algorithm identifies first the set of states (different from $\bar{\bf x}=\delta_{N}^{j}$) that have $\bar{\bf x}=\delta_{N}^{j}$ as their successor, namely ${\mathcal{S}}_{1}$. For each such state $\delta^{r}_{N}$, it memorizes in ${\bf u}_{r}^{\bar{\bf x}}$ one of the values of the input that allows the transition from $\delta^{r}_{N}$ to $\bar{\bf x}=\delta^{j}_{N}$. Then it identifies the set ${\mathcal{S}}_{2}$ of the states that do not belong to ${\mathcal{S}}_{0}\cup{\mathcal{S}}_{1}=\{\bar{\bf x}\}\cup{\mathcal{S}}_{1}$, but have a successor in ${\mathcal{S}}_{1}$. Again, for each $\delta^{r}_{N}\in{\mathcal{S}}_{2}$, it memorizes one of the values of the input that allows the transition from $\delta^{r}_{N}$ to some state in ${\mathcal{S}}_{1}$. Since the set of states of a BCN is finite, there exists $d\geq 1$ such that ${\mathcal{S}}_{d}=\emptyset$. If at that stage $\mathcal{S}(\bar{\bf x})=\bigcup_{i=0}^{d-1}\mathcal{S}_{i}=\mathcal{L}_{N}$, this means that $\bar{\bf x}=\delta^{j}_{N}$ is globally reachable and $\mathcal{S}(\bar{\bf x})$ is its basin of attraction.
We are now ready to state Lemma 9, whose proof is straightforward from the previous analysis.
Lemma 9.
Given $\bar{\bf x}\in\mathcal{L}_{N}$, the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for reachability of $\bar{\bf x}$ if and only if Algorithm 2 terminates with an affermative answer.
By putting together the results obtained so far, and leveraging the model-based results in Proposition 4 and subsequent comments, we can provide the solution to Problem 1, namely a characterization of informativity for stabilization by state feedback.
Theorem 10.
Problem 1 is solvable (i.e., the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for stabilization of ${\bf x}_{e}$ by state feedback) if and only if
  • (i)
    ${\bf x}_{e}\in\mathcal{X}_{e}^{d}$;
  • (ii)
    the data $({\bf x}_{d}$, ${\bf u}_{d})$ are informative for reachability of ${\bf x}_{e}$.
Moreover, the $r$-th column of a possible feedback matrix $K\in{\mathcal{L}}_{M\times N}$ is given by  ${\rm col}_{r}(K)={\bf u}_{r}^{{\bf x}_{e}},\forall r\in[1,N],$ where ${\bf u}_{r}^{{\bf x}_{e}}$ is the logical vector determined in Algorithm 2.
Proof.
(Necessity). Assume, by contradiction, that condition (i) does not hold, namely ${\bf x}_{e}\notin\mathcal{X}_{e}^{d}$. This means that there exists at least one BCN in $\mathcal{B}_{d}$ for which ${\bf x}_{e}$ is not an equilibrium point and hence, by Proposition 4, such network cannot be stabilized to ${\bf x}_{e}$. This contradicts the fact that Problem 1 is solvable. Assume now that ${\bf x}_{e}\in\mathcal{X}_{e}^{d}$, but condition (ii) fails to hold. This means that there exists at least one BCN in $\mathcal{B}_{d}$ for which ${\bf x}_{e}$ is not globally reachable. One of such BCNs is simply obtained by imposing that all the transitions that cannot be deduced from data are self-loops. By Proposition 4, such BCNs cannot be stabilized, and hence again Problem 1 is not solvable.
(Sufficiency). It follows directly from Proposition 4 and previous comments and results.
The final statement follows from the fact that, for every state $\delta_{N}^{r}$ of every BCN in ${\mathcal{B}}_{d}$, ${\bf u}_{r}^{{\bf x}_{e}}$ is exactly the/an input that, starting from this state, ensures the correct transition towards reaching ${\bf x}_{e}$, by construction (see comments after Algorithm 2). ∎
Remark 11.
A necessary condition for the data $({\bf x}_{d}$, ${\bf u}_{d})$ to be informative for stabilization to ${\bf x}_{e}\in\mathcal{L}_{N}$ by state feedback is that the matrix $X_{p}$ is of full row rank [^2] This imposes $T\geq N$., or, equivalently, each state of $\mathcal{L}_{N}$ appears at least once as a column of $X_{p}$. Indeed, in order to check that the state ${\bf x}_{e}=\delta^{j}_{N}$ belongs to $\mathcal{X}_{e}^{d}$, namely is an equilibrium point of all BCNs compatible with the data, there must be a column of $X_{p}$ that coincides with ${\bf x}_{e}$ (and whose corresponding column in $X_{f}$ also coincides with ${\bf x}_{e}$). On the other hand, if there is some state, say $\delta_{N}^{q}$, $q\neq j$, that does not appear as a column of $X_{p}$, we cannot deduce from the data its successor for any $i$-th input, $i\in[1,M]$. This implies that in all the BCNs compatible with the data, the successors of $\delta_{N}^{q}$ can be arbitrarily chosen. As a consequence, in all the BCNs in $\mathcal{B}_{d}$ for which ${\rm col}_{q}(\tilde{L}_{i})=\delta_{N}^{q},\forall i\in[1,M]$, ${\bf x}_{e}$ would not be globally reachable. This prevents the data $({\bf x}_{d}$, ${\bf u}_{d})$ to be informative for reachability of ${\bf x}_{e}$.
4 Data-driven safe control
In Fornasini & Valcher (2016) it has been provided a complete characterization of the problem solvability.
Proposition 12.
  • (i)
  • (ii)
Problem 2.
When Problem 2 is solvable, the data $({\bf x}_{d}$, ${\bf u}_{d})$ will be called informative for safe control (with respect to the unsafe set ${\mathcal{X}}_{u}$). The problem solution relies on the tools developed to solve Problem 1. First of all, we need to identify from data if each state in ${\mathcal{X}}_{s}$ has (at least) one successor in ${\mathcal{X}}_{s}$. This amounts to verifying if for every $\delta^{j}_{N}\in{\mathcal{X}}_{s}$ there exists $k\in[1,T]$ and $\delta^{\ell}_{N}\in{\mathcal{X}}_{s}$ such that
$\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{\ell}\end{bmatrix}.$ (15)
If this is the case we can set
${\bf u}_{j}^{*}\triangleq U_{p}\delta_{T}^{k}.$ (16)
If the previous check gives a positive answer, then condition (i) in Proposition 12 holds for every BCN in ${\mathcal{B}}_{d}$. At this point we need to check that condition (ii) holds for every BCN in ${\mathcal{B}}_{d}$. This is possible by extending Algorithm 2.
Algorithm 3: Reachability of ${\mathcal{X}}_{s}$.
@@ 4 Data-driven safe control
           - The set ${\mathcal{X}}_{s}=\{\delta^{j_{1}}_{N},\delta^{j_{2}}_{N},\dots,\delta^{j_{d}}_{N}\}$.
Output: - Is ${\mathcal{X}}_{s}$ globally reachable for every BCN in ${\mathcal{B}}_{d}$?
               Yes/No.
             - The basin of attraction ${\mathcal{S}}^{*}\triangleq\bigcup_{{\bf x}\in{\mathcal{X}}_{s}}\mathcal{S}({\bf x})$ of
               ${\mathcal{X}}_{s}$.
@@
for $r\in[1,N]$, do
     if $\delta^{r}_{N}\in\mathcal{X}_{s}$, then
         ${\bf u}_{r}\leftarrow{\bf u}_{r}^{*}$ (see (16))
     end if
end for
Set $d=0$, ${\mathcal{S}}_{0}=\mathcal{X}_{s}$.
@@
else set $\mathcal{S}^{*}=\bigcup_{i=0}^{d-1}\mathcal{S}_{i}$ and go to the Conclusion

### v2

@@ 4 Reachability property
and the associated digraph $\mathcal{D}(L_{tot})\triangleq(\mathcal{V},\mathcal{E}_{tot})$. The state ${\bf x}_{f}=\delta_{N}^{i}$ is reachable from ${\bf x}_{0}=\delta_{N}^{j}$ if and only if there exists $\tau\in\mathbb{Z}_{+}$ such that $L_{tot}^{\tau}]_{ij}=1$, or equivalently, there is a path of length $\tau$ from $j$ to $i$ in $\mathcal{D}(L_{tot})$. [^3] In the sequel, we will interchangeably write that a node (in the digraph) or a state (of the BCN) is reachable from another node or state. More generally, as there is a bijective correspondence between states and nodes, with a slight abuse of terminology, we will often refer to state $\delta^{i}_{N}$ to denote node $i$ in the digraph, and vice versa. This is the case if and only if $(\delta^{i}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}L_{tot}^{\tau}\right)\delta^{j}_{N}>0$. Consequently, the state ${\bf x}_{f}=\delta^{i}_{N}$ (the set ${\mathcal{X}}=\{\delta^{i_{1}}_{N},\delta^{i_{2}}_{N},\dots,\delta^{i_{d}}_{N}\}$) is globally reachable if and only if all entries of the row vector $(\delta^{i}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}L_{tot}^{\tau}\right)$ (of the row vector $(\sum_{\ell=1}^{d}\delta^{i_{\ell}}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}L_{tot}^{\tau}\right)$) are positive. This implies that a BCN is reachable if and only if the matrix $\sum_{\tau=0}^{N-1}L_{tot}^{\tau}$ has all positive entries, which means that $L_{tot}$ is irreducible, or equivalently $\mathcal{D}(L_{tot})$ is strongly connected.
We are now ready to introduce the notion of informativity for reachability of a set of states $\mathcal{X}$ that extends Definition 3.7 in Li et al. (2025).
@@
Given a set ${\mathcal{X}}\subset\mathcal{L}_{N}$ (in particular, a state ${\bf x}_{f}\in{\mathcal{L}}_{N}$), we say that the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for reachability of $\mathcal{X}$ if the set $\mathcal{X}$ (the state ${\bf x}_{f}$) is globally reachable for all the BCNs in ${\mathcal{B}}_{d}$.
Since the reachability of some state ${\bf x}_{f}$ from another state ${\bf x}_{0}$ is related to the existence of a finite number of state transitions that lead the state from ${\bf x}_{0}$ to ${\bf x}_{f}$, it is clear that when we try to verify this property for all BCNs in ${\mathcal{B}}_{d}$, we can rely only on the state transitions that are revealed by the data. This can be performed by first identifying the best possible approximation of $L_{tot}$ based on data, say $L_{tot}^{d}\in{\mathcal{B}}^{N\times N}$, and then verifying the reachability properties as explained in the model-based approach by referring to such matrix. Note that in general (differently from $L_{tot}$) the matrix $L_{tot}^{d}$ may have zero columns. This happens if there exists $j\in[1,N]$ such that $\delta^{j}_{N}$ does not appear as a column of $X_{p}$. We have the following result.
Proposition 6.
Set $L_{tot}^{d}\triangleq X_{f}\odot_{\mathcal{B}}X_{p}^{\top}$. Then for every $(i,j)\in[1,N]\times[1,N]$, there exists $k\in[1,T]$ such that
$\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{i}\end{bmatrix}$, (16)
if and only if $L_{tot}^{d}]_{ij}=1$. Therefore the data are informative for reachability of the set ${\mathcal{X}}=\{\delta^{i_{1}}_{N},\delta^{i_{2}}_{N},\dots,\delta^{i_{d}}_{N}\}$ (of the state ${\bf x}_{f}=\delta^{i}_{N}$) if and only if all entries of the row vector $(\sum_{\ell=1}^{d}\delta^{i_{\ell}}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}(L_{tot}^{d})^{\tau}\right)$ (of the row vector $(\delta^{i}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}(L_{tot}^{d})^{\tau}\right)$) are positive.
Proof.
The first statement follows from Lemma 3.8 in Li et al. (2025). The second part is immediate from the first statement and the comments before Proposition 6. ∎
Proposition 6 provides an extension of Theorem 3.9 in Li et al. (2025), where a characterization of informativity for (the overall) reachability is provided. Indeed, differently from Li et al. (2025), we focus on the reachability of a set or a state, since this will be useful in the following sections. Moreover, as an interesting byproduct, we have deduced the matrix $L_{tot}^{d}$ that represents the best possible “under-approximation" of the matrix $L_{tot}$, in the sense that it is the largest Boolean matrix that satisfies $L_{tot}^{d}\leq L_{tot}$ for all the matrices $L_{tot}$ of the BCNs in ${\mathcal{B}}_{d}$.
It is worth underlining that Proposition 6 only provides an answer to the question of whether or not data are informative for reachability, but it does not offer any insight into the way one can select an input sequence that achieves the goal. To this end, we introduce an algorithm, which also allows us to explore another concept that will be useful in the subsequent analysis.
@@
Given a BCN described as in (6a), the basin of attraction ${\mathcal{S}}({\bf x}_{f})$ of a state ${\bf x}_{f}\in{\mathcal{L}}_{N}$ is the set of states ${\bf x}_{0}\in{\mathcal{L}}_{N}$ from which it is possible to reach ${\bf x}_{f}$.
Clearly, a state ${\bf x}_{f}$ is globally reachable if and only if its basin of attraction ${\mathcal{S}}({\bf x}_{f})$ coincides with ${\mathcal{L}}_{N}$.
Algorithm 1, below, provides a way to determine the basin of attraction $\mathcal{S}$$^{*}$ of a set $\mathcal{X}$}, by this meaning the union of the basins of attractions of the states in $\mathcal{X}$}. For each state in ${\mathcal{S}}^{*}\setminus{\mathcal{X}}$ the algorithm also returns a possible choice for the input that if applied to that state at a certain time instant $\bar{t}$, it ensures that at the next time instant $\bar{t}+1$ the distance between the state and the target set $\mathcal{X}$ decreases. Therefore, it readily follows that, by applying the inputs provided by the algorithm, if the set $\mathcal{X}$ is globally reachable, it will be reached in a finite number of steps.
Algorithm 1 Reachability of $\mathcal{X}$}
@@
     - The set ${\mathcal{X}}=\{\delta^{i_{1}}_{N},\delta^{i_{2}}_{N},\dots,\delta^{i_{d}}_{N}\}$.
Output: - A possible choice of inputs $({\bf u}_{1},{\bf u}_{2},\dots,{\bf u}_{N})$;      - The basin of attraction ${\mathcal{S}}^{*}\triangleq\bigcup_{{\bf x}\in{\mathcal{X}}}\mathcal{S}({\bf x})$ of
      $\mathcal{X}$}.
@@
Set $d=0$, ${\mathcal{S}}_{0}=\mathcal{X}$.
@@
else set $\mathcal{S}^{*}=\bigcup_{i=0}^{d-1}\mathcal{S}_{i}$